OpenStudy (anonymous):

Solve negative 7 over 3, the whole multiplied by x minus 3 equals- 52. halp

OpenStudy (anonymous):

Is it like \(\Large \frac{7}{3}x - 3= -52\)?

OpenStudy (anonymous):

yes Not sure if that is pi though

OpenStudy (anonymous):

-7/3x -3 = -52

OpenStudy (anonymous):

pi? ?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

-7/3x -3 = -52

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

1. transpose -3 to the right side by adding both sides by 3 then simplify 2. multiply both sides by 3 then simplify 3. divide both sides by -7 and simplify

OpenStudy (anonymous):

i think 21

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

please disregard step 3

OpenStudy (anonymous):

2. multiply both sides by 3x then simplify 3. divide both sides by the coefficient of x